OpenStudy (anonymous):

Find the volume below the surface z=(sqrt(16-x^2-y^2)) which lies above the cone Z^2=3x^2+3y^2

OpenStudy (anonymous):

First, set up your integral.

OpenStudy (anonymous):

But before that shouldnt we turn the variables into polar?

OpenStudy (anonymous):

I was thinking that was part of the setting up process. Go for it.

OpenStudy (anonymous):

haha ok sorry. Lemme see....

OpenStudy (anonymous):

z=sqrt(16-(rcostheta)^2-(rsintheta)^2 =sqrt(16-r^2)

OpenStudy (anonymous):

z=-r+4

OpenStudy (anonymous):

Good so far.

OpenStudy (anonymous):

z^2=3(x^2+y^2) z^2=3(r^2)

OpenStudy (anonymous):

Substitute for z.

OpenStudy (anonymous):

r=sqrt(8)

OpenStudy (anonymous):

Hm, that's not what I got. Let me recheck.

OpenStudy (anonymous):

Yeah, I didn't get sqrt(8). Check your math. One of us is wrong.

OpenStudy (anonymous):

Wait what did you get?

OpenStudy (anonymous):

i got r=2

OpenStudy (anonymous):

@Mertsj please help